OpenStudy (belle):

3/8 of the no of english books in a library is equal to 1/9 of the no of chinese books. There are 2 100 chinese and english books altogether. How many more Chinese than English books are there in the library?

OpenStudy (anonymous):

3/8x = 1/9(2100-x) 72*3 x = 72*1(2100-x) ---- ------------ 8 9 9 * 3x = 8 * 1*(2100-x) 27x = 8(2100 - x) 27x = 16800 - 8x 35x = 16800 x = 480 English Books 2100 - 480 = 1620 Chinese Books

OpenStudy (belle):

thanks Blexting. you are great.